a)     Suppose that UNZA is a country with a population of 321,166,640 people of which the number of unemployed people is 121,166,640. The employment rate in this economy is 0.104, and the labour force participation rate 72.5%.


1
Expert's answer
2021-07-08T15:20:54-0400

(i) labor force =unemployed÷unemployment rate

121,166,640÷0.104×100=116,505,384,600

(ii) unemployed people =121,166,640

(iii) working age =employed÷ labor force participation rate

11,638,521,800÷0.0725=160,531,335,200

(I)GDP per Capita =output per person÷country population

625,234,567÷321,166,640=1.94676